OpenStudy (anonymous):

If the radius of a circle is 8, what is the circumference of the circle.

OpenStudy (anonymous):

Radius is always half the circumference.

OpenStudy (anonymous):

16

OpenStudy (anonymous):

\[C = 2πr\]

OpenStudy (anonymous):

Plug in 8 for r. \[C = 2(π)(8)\]\[C = ?

OpenStudy (anonymous):

C=2*22/7*8 =16*22/7 =50.286

OpenStudy (anonymous):